A farmer exerts a force upon a 3.29 kg pile of soil to both lift it and set into motion. The soil leaves the shovel with a speed of 2.94 m/s at a height of 0.562 m. Determine the work done upon the pile of snow.



1
Expert's answer
2021-06-02T09:34:58-0400

Kei + Pei + Wapplied = Kef + Pef

Initially, snow was at rest on the ground:

PEin=KEin=0

When teacher exerts force on snow, it does some work upon snow, which is equal to kinetic energy + potential energy gained by snow.

Work=PEfi+ KEfi

"W = mgh + \frac{1}{2}mV^2 \\

= 3.29 \times 9.8 \times 0.562 + \frac{1}{2} \times 3.29 \times 2.94^2 = 32.34 \;J"

Answer: 32.34 J